OpenStudy (anonymous):

solve please 3x + 2/6y=6 7x - 2y=13

OpenStudy (anonymous):

ok this is substitution. so have y or x by itself. like y= or x= (3x+2)/6y=6 3x+2 =36y (3x/36) + (2/36) = y (<-- see what I mean^) now plug y into the sec formula. 7x-2[(3x/36) + (2/36)] = 13 <-- so just multiply everything out and solve for x. FYI it doesn't matter which equation you start first. when getting y= or x= (hope I make some sense to you)

OpenStudy (anonymous):

yeah ..it make some sense

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

thanks